Hello! this is an s.o.s. fI need to figure out this problem for class and keep getting stuck. I am supose to translate this problem into two a pair of linear equations in two variables. solve the equation using the elimination or substitute method.
Roberto invested some money at 7%, and then invested $2000 more than twice this amount at 9%. His total annual income from the two investments was $2930. How much was invested at 9% ?
I started out like this, but keep getting stumped!!!
x(0.07) + x*2+2000=2930 or
2930 - s(0.07)=n
2930- 2s+2000(0.09)=n
See wwhat I mean?


Answer by nerdybill(7384)

Roberto invested some money at 7%, and then invested $2000 more than twice this amount at 9%. His total annual income from the two investments was $2930. How much was invested at 9% ?
Let x = amount invested at 7%
then
2x+2000 = amount invested at 9%
.
x(0.07) + (2x+2000)(0.09) = 2930
.07x + 0.18x+180 = 2930
0.25x+180 = 2930
0.25x = 2750
x = $11000 (amount invested at 7%)
.
Amount invested at 9% then:
2x+2000 = 2(11000)+2000 = 22000+2000 = $24,000 (amount invested at 9%)